OpenStudy (anonymous):

Brad walks and jogs to schooll every day. He averages 5 km/hr walking and 9 km/hr jogging. The distane from home to shool is 6 km, and the total trip takes 1 hour. How far does Brad walk?

OpenStudy (anonymous):

6-x/9 + x/5 = 1 9x +30 - 5x = 45 4x = 15 x = 3.75 km He walks 3.75 km

OpenStudy (anonymous):

Let x be the time taken he walk to school every day and y be the time taken he run to school every day. Use x and y to form 2 distinct equations to solve for the answer. Given that total trip takes 1 hour, x+y = 1 - (1) Give the speed of 5km/hr for walking and 9km/hr for jogging, 5x+9y = 6 -(2) Taking eqn.(1) and multiply by 5. x+y= 1 5x+5y = 5 - (3) take eqn.2 and subtract eqn.3 5x+9y - (5x +5y) = 6-5 4y = 1 y= 1/4 Since y=1/4 , from eqn.1 x+ (1/4) = 1 x = 3/4 He walks for 3/4 hr = 45mins.